ABC Ltd. uses EOQ logic to determine the order quantity for its various components and is planning its orders. The Annual consumption is 80,000 units, Cost to place one order is Rs. 1,200, Cost per unit is Rs. 50 and carrying cost is 6% of Unit cost. Find EOQ, No. of order per year, Ordering Cost and Carrying Cost and Total Cost of Inventory.


1
Expert's answer
2021-09-01T11:51:53-0400

"EOQ=\sqrt{(\frac{2 \times RU \times OC }{UC \times CC\%})}"

Where,

EOQ is the Economic Order Quantity.

RU is the Required Units

CC is the Carrying Cost

OC is Ordering Cost for a Unit.

UC is the Inventory Unit Cost.

a) EOQ (Economic Order Quantity)

EOQ = "\sqrt{(\frac{2 \times 80000\times1200}{50 \times 6\%})}"

"= 8000\ units"


b) No. of order per year

"No. \ of \ per \ year \ = \frac{Annual Requirements } {Economic\ Order\ Quantity}"

"=\frac{80000}{8000}"

"=\ 10"

c) Ordering Cost

Ordering Cost = No. of order per year x fixed Ordering Cost

"= 1200 \times \ 10"

"= 12000\ Rupees"

d) Carrying Cost (CC)

"CC = Carrying\ Cost\% \times \ Cost \ per \ unit \times \frac{EOQ}{2}"

"=50 \times 0.006 \times \frac{8000}{2}"

"= 12000\ Rupees"

e) Total Cost of Inventory

"Total\ Cost\ of \ Inventory\ = Ordering\ Cost+ Carrying\ Cost"

"= 12000 + 12000"

"=24000 \ Rupees"